# SERA-H: Beyond Native Sentinel Spatial Limits for High-Resolution Canopy Height Mapping

### v2

@@ Abstract
High-resolution mapping of canopy height is essential for forest management and biodiversity monitoring. Although recent studies have led to the advent of deep learning methods using satellite imagery to predict height maps, these approaches often face a trade-off between data accessibility and spatial resolution. To overcome these limitations, we present SERA-H, an end-to-end model combining a super-resolution module (EDSR) and temporal attention encoding (UTAE). Trained under the supervision of high-density LiDAR data (ALS), our model generates 2.5 m resolution height maps from freely available Sentinel-1 and Sentinel-2 (10 m) time series data. Evaluated on an open-source benchmark dataset in France, SERA-H, with a MAE of 2.6 m and R2 of 0.82, not only outperforms standard Sentinel-1/2 baselines but also achieves performance comparable to or better than methods relying on commercial very high-resolution imagery (SPOT-6/7, PlanetScope, Maxar). These results demonstrate that combining high-resolution supervision with the spatiotemporal information embedded in time series enables the reconstruction of details beyond the input sensors’ native resolution. SERA-H opens the possibility of freely mapping forests with high revisit frequency, achieving accuracy comparable to that of costly commercial imagery.
@@ 1 Introduction
Canopy height maps can be derived by processing data from Airborne Laser Scanning (ALS) and NASA’s Global Ecosystem Dynamics Investigation (GEDI). Direct measurements are also obtained from National Forest Inventories (NFIs) at point locations. While the GEDI mission has established itself as the reference for monitoring the vertical structure of forests on a global scale, it operates by nature as a sampling mission providing sparse spatial and temporal samples (so-called footprints). ALS is the reference for high-resolution local canopy height mapping, but its deployment remains costly and geographically limited. As for NFI inventories, although they provide direct in-situ measurements, they remain limited to a small number of plot locations and do not allow for continuous spatial monitoring. To overcome the discontinuity of these reference measurements, deep learning approaches have emerged as an effective solution. These approaches exploit the continuous spatial and temporal coverage of optical (e.g., Sentinel-2, Landsat, SPOT) and radar (e.g., Sentinel-1, PALSAR-2) imagery to predict canopy height maps under the supervision of LiDAR acquisitions, such as those from GEDI or ALS data. This enables large-scale canopy height prediction using data that is both frequently updated and cost-effective.
@@
Recently, open access to ALS data has expanded significantly (LiDAR HD program  [^1] https://geoservices.ign.fr/lidarhd, Fischer et al. (2025)) providing extensive coverage of high-quality data. Beyond serving as the most accurate reference for locally estimating canopy height (Alexander et al., 2018; Lefsky et al., 2005), ALS datasets offer a very high spatial resolution (typically 1-2 m). On the other hand, the highest resolution provided by open-access satellite data comes from the Sentinel missions, part of ESA’s Copernicus Earth Observation program. At 10 m, the resolution of Sentinel-1 and Sentinel-2 satellite images remains considerably coarser than the level of detail offered by ALS-based canopy height maps. To match the resolution of the input imagery with the reference data for model training, the most common approach is to downgrade the resolution of the canopy height derived from ALS data to align it with the Sentinel resolution (Liu et al., 2023; Su et al., 2025b, a), inevitably resulting in significant information loss. Here, we propose an alternative approach: to increase the resolution of Sentinel images to match the fidelity of ALS-derived canopy height maps.
@@
At the same time, a recent study has demonstrated the value of directly exploiting the time series of raw Sentinel-2 images, in contrast to the standard practice of using annual or seasonal mosaics, to estimate canopy height (Pauls et al., 2025). By avoiding the loss of information inherent in annual compositing, this model exploits the seasonal and intra-annual dynamics of input Sentinel-2 images, significantly improving the accuracy of height predictions, particularly for tall trees. Another study by Sirko et al. (2023) combined a super-resolution approach and multi-temporal images, by simultaneously using spatial and temporal variations in successive acquisitions to produce higher-resolution results in various urban remote sensing contexts. Their model has proven effective for tasks such as building segmentation, centroid and height estimation, and road segmentation. Fundamentally, this performance relies on exploiting the sub-pixel shifts inherent in revisited acquisitions, which allows the model to resolve aliasing and reconstruct fine spatial details by leveraging this temporal redundancy. However, to our knowledge, no study combining super-resolution and multi-temporal images for forestry applications exists to date.
@@ 2.1 Data
The proposed approach aims to transform Sentinel-1 and Sentinel-2 time series (10 m) into high-resolution canopy height maps (2.5 m) using super-resolution and temporal regression. This section details the datasets used for training and evaluation. We first describe the Open-Canopy benchmark (Section˜2.2), which provides the ALS-based reference height maps, the dataset splits, and the vegetation mask. Then, we present the acquisition and processing of the input Sentinel satellite time series (Section˜2.2.1).
@@ 2.2 Open-Canopy Dataset: Reference Data, Splits, and Vegetation Mask
To train and evaluate our model, we used the Open-Canopy dataset (Fogel et al., 2025), a public benchmark specifically designed for high-resolution canopy height estimation. Derived from the French LiDAR HD campaigns acquired between 2021 and 2023222https://geoservices.ign.fr/lidarhd, the Open-Canopy dataset provides ALS-based canopy height maps at 1.5 m resolution which served as reference for height prediction. For the purpose of our study, we downsampled these reference images to 2.5 m using max-pooling aggregation.
The dataset spans 87,383 patches of $1\times 1$ km distributed across France, ensuring a wide diversity of forest types and topographies. We adhered to the fixed spatial partition defined by the Open-Canopy benchmark: 66,339 km2 for training, 7,369 km2 for validation, and 13,675 km2 for testing (Fig.˜1), with a 1 km buffer systematically applied around test areas to prevent spatial leakage.
Finally, to ensure that performance metrics focused on relevant areas, we restricted the evaluation to vegetation-covered zones using the vegetation mask provided by the Open-Canopy dataset. This mask was constructed by the union of a LiDAR-derived mask (height >1.5>1.5 m) and IGN forest polygons [^3] https://geoservices.ign.fr/bdforet. Note that this mask was exclusively applied during the validation and test phases for metric calculation and was not used during training.
Figure 1: Spatial splits from the Open-Canopy dataset. The inset zooms into a representative region to show the spatial arrangement of the different training, test, and validation tiles. A 1 km buffer zone is applied around the test areas to avoid any spatial leakage.
@@ 2.2.1 Satellite Data: Sentinel-1 and Sentinel-2
Unlike the Open-Canopy benchmark which relied on commercial SPOT-6/7 imagery, our model deployed input imagery from the Sentinel-1 and Sentinel-2 sensor systems, key components of the European Space Agency’s (ESA) Copernicus Earth observation program. These satellite systems provide complementary radar and optical data with global coverage and high temporal resolution. We retrieved these images using the geefetch package [^4] https://geefetch.readthedocs.io/en/latest/ (Belouze et al., 2025), which interfaces with the Google Earth Engine (GEE) API [^5] https://earthengine.google.com/. For each patch of the Open-Canopy Dataset, all Sentinel images available within a 120-day window centered on the acquisition date of the corresponding ALS data were retrieved. To minimize the influence of atmospheric perturbations, Sentinel-2 data were subjected to a cloud-masking procedure. Images with more than 40% cloudy pixels or 20% high-probability cloud coverage were excluded. Individual pixels with a cloud probability above 40% or flagged in Sentinel-2’s QA60 band were masked.
For each raw Sentinel-2 ($S2$) image, the closest raw Sentinel-1 ascending ($S1_{asc}$) and descending ($S1_{dsc}$) images in time were associated to form a triplet ($S2$, $S1_{asc}$, $S1_{dsc}$). Ascending and descending acquisitions were distinguished because they have opposite observation angles, offering two complementary viewpoints. Furthermore, combining Sentinel-2 optical images and Sentinel-1 radar images allowed simultaneous exploitation of the spectral and structural properties of the canopy. Across the dataset, we obtained an average of 13.51 triplets per patch.
For Sentinel-2, we selected 10 spectral bands that are relevant for vegetation characterization: the visible RGB channels (B2, B3, B4), the vegetation red-edge bands (B5, B6, B7), the near-infrared (NIR) bands (B8, B8A), and the short-wave infrared (SWIR) bands (B11, B12). Each Sentinel-1 image contained 2 bands for VV and VH polarizations for both ascending and descending orbits. Consequently, each input triplet comprised a total of 14 channels (10 optical and 4 radar), all harmonized at a resolution of 10 m using nearest-neighbor resampling.
@@
Figure 2: Examples of input and reference data used in SERA-H. (a) Sentinel-2 optical image, (b) Sentinel-1 ascending radar image, and (c) Sentinel-1 descending radar image together form one of the triplet of the time series used as model input. (d) The corresponding ALS-derived canopy height map, resampled at 2.5 m, serves as reference.
@@ 2.3 Model
Figure 3: Overview of the SERA-H model workflow. Each input sample consists of a temporal sequence of triplets, where each triplet is composed of one Sentinel-2 optical image and its temporally closest Sentinel-1 ascending and descending acquisitions ($S2$, $S1_{asc}$, $S1_{dsc}$). This yields an input tensor of size $T\times C\times H\times W$, where $T$ is the fixed number of triplets in the temporal sequence, $C$ the total number of spectral and radar channels per triplet, and $H$ and $W$ the spatial dimensions of each image. The sequence is first processed by the EDSR super-resolution module, which upsamples all channels by $\times 4$ to match the reference ALS resolution. The super-resolved images are then passed through the UTAE encoder-decoder, which compresses the full temporal sequence using a temporal attention mechanism and outputs a single high-resolution canopy height map ($1\times 1\times 4H\times 4W$). Training is performed end-to-end using a Smooth L1 loss against ALS reference data, with gradients propagating through both UTAE and EDSR.
@@ 2.3.1 Enhanced Deep Super-resolution (EDSR) Module
Super-resolution reconstructs a high-resolution image from low-resolution inputs by recovering missing details to improve spatial clarity and pixel density. In the field of super-resolution models, Generative Adversarial Networks (GANs) are widely used (Ledig et al., 2016; Jiang et al., 2019; Guo et al., 2022). These models are based on competition between a generator, which synthesizes plausible high-resolution images, and a discriminator responsible for distinguishing between real and generated images. Although this competitive training promotes high visual realism, it encourages the generator to prioritize perceptual plausibility at the expense of strict fidelity to reality, often resulting in the appearance of unwanted artifacts (so-called hallucinations). Consequently, for our task, adversarial approaches were excluded.
Instead, we opted for the Enhanced Deep Super-Resolution (EDSR) model (Lim et al., 2017), which has demonstrated state-of-the-art performance in upsampling satellite imagery while being less prone to artifacts (Karwowska and Wierzbicki, 2022; Ren et al., 2023). Derived from the ResNet architecture, EDSR optimizes the model by removing unnecessary normalization layers. This simplification allows for a much larger and more powerful network to be built. We thus adopted this architecture in our workflow to upsample each individual image within the time series, increasing the spatial resolution by a factor of 4.
@@ 2.4.1 Training Data Sampling
During training, a $1\times 1$ km tile was randomly sampled from the Open-Canopy dataset. To construct the time series, we identified the triplets ($S2,S1_{asc},S1_{dsc}$) temporally closest to the reference ALS acquisition of this zone. The date assigned to each triplet corresponded to the Sentinel-2 acquisition date. Tiles containing fewer than 4 distinct triplets were discarded from the training set. For the remaining tiles, as the model requires a fixed sequence length of $T=16$, we padded sequences with an insufficient number of triplets by randomly duplicating existing ones or selected the 16 closest ones if more are available. These images were subsequently cropped into $64\times 64$ pixel patches to match the model’s input dimensions. Corresponding ALS canopy height references were extracted at the same coordinates with a 2.5 m resolution, yielding $256\times 256$ pixel reference images.

### v3

@@ Abstract
High-resolution mapping of canopy height is essential for forest management and biodiversity monitoring. Although recent studies have led to the advent of deep learning methods using satellite imagery to predict height maps, these approaches often face a trade-off between data accessibility and spatial resolution. To overcome these limitations, we present SERA-H, an end-to-end model combining a super-resolution module (EDSR) and temporal attention encoding (UTAE). Trained under the supervision of high-density LiDAR-derived Canopy Height Models (CHM), our model generates 2.5 m resolution height maps from freely available Sentinel-1 and Sentinel-2 (10 m) time series data. Evaluated on an open-source benchmark dataset in France, SERA-H, with a MAE of 2.6 m and R2 of 0.82, not only outperforms standard Sentinel-1/2 baselines but also achieves performance comparable to or better than methods relying on commercial very high-resolution imagery (SPOT-6/7, PlanetScope, Maxar). These results demonstrate that combining high-resolution supervision with the spatiotemporal information embedded in time series enables the reconstruction of details beyond the input sensors’ native resolution. SERA-H opens the possibility of freely mapping forests with high revisit frequency, achieving accuracy comparable to that of costly commercial imagery.
@@ 1 Introduction
Canopy height maps can be derived by processing data from Airborne Laser Scanning (ALS) and NASA’s Global Ecosystem Dynamics Investigation (GEDI). Direct measurements are also obtained from National Forest Inventories (NFIs) at point locations. While the GEDI mission has established itself as the reference for monitoring the vertical structure of forests on a global scale, it operates by nature as a sampling mission providing sparse spatial and temporal samples (so-called footprints). ALS is the reference for high-resolution local canopy height mapping, but its deployment remains costly and geographically limited. As for NFI inventories, although they provide direct in-situ measurements, they remain limited to a small number of plot locations and do not allow for continuous spatial monitoring. To overcome the discontinuity of these reference measurements, deep learning approaches have emerged as an effective solution. These approaches exploit the continuous spatial and temporal coverage of optical (e.g., Sentinel-2, Landsat, SPOT) and radar (e.g., Sentinel-1, PALSAR-2) imagery to predict canopy height maps under the supervision of LiDAR acquisitions, such as those from GEDI or ALS data. To achieve this, models leverage optical reflectance, which correlates with vertical forest structure (Woodcock et al., 1994), and radar backscatter, which depends on the geometric properties of the forest. However, these methods encounter saturation effects specific to each type of signal, which can limit their ability to make accurate predictions about large trees. Despite these constraints, this framework enables large-scale canopy height prediction using data that is both frequently updated and cost-effective.
@@
Recently, open access to ALS data has expanded significantly (LiDAR HD program  [^1] https://geoservices.ign.fr/lidarhd, Fischer et al. (2025)) providing extensive coverage of high-quality data. Beyond serving as the most accurate reference for locally estimating canopy height (Alexander et al., 2018; Lefsky et al., 2005), ALS datasets offer a very high spatial resolution (typically 1-2 m). On the other hand, the highest resolution provided by open-access satellite data comes from the Sentinel missions, part of ESA’s Copernicus Earth Observation program. At 10 m, the resolution of Sentinel-1 and Sentinel-2 satellite images remains considerably coarser than the level of detail offered by ALS-based canopy height maps. To match the resolution of the input imagery with the reference data for model training, the most common approach is to downgrade the resolution of the canopy height derived from ALS data to align it with he satellite’s native resolution. (Liu et al., 2023; Su et al., 2025b, a), inevitably resulting in significant information loss. Here, we propose an alternative approach: to increase the resolution of Sentinel images to match the fidelity of ALS-derived canopy height maps.
@@
At the same time, a recent study has demonstrated the value of directly exploiting the time series of raw Sentinel-2 images, in contrast to the standard practice of using annual or seasonal mosaics, to estimate canopy height (Pauls et al., 2025). By avoiding the loss of information inherent in annual compositing, this model exploits the seasonal and intra-annual dynamics of input Sentinel-2 images, significantly improving the accuracy of height predictions, particularly for tall trees. Another study by Sirko et al. (2023) combined a super-resolution approach and multi-temporal images, by simultaneously using spatial and temporal variations in successive acquisitions to produce higher-resolution results in various urban remote sensing contexts. This model has proven effective for tasks such as building segmentation, centroid and height estimation, and road segmentation. Fundamentally, this performance relies on exploiting the sub-pixel shifts inherent in revisited acquisitions, which allows the model to resolve aliasing and reconstruct fine spatial details by leveraging this temporal redundancy. However, to our knowledge, no study combining super-resolution and multi-temporal images for forestry applications exists to date.
@@ 2.1 Data
The proposed approach aims to transform Sentinel-1 and Sentinel-2 time series (10 m) into high-resolution canopy height maps (2.5 m) using super-resolution and temporal regression. This section details the datasets used for training and evaluation. We first describe the Open-Canopy benchmark (Section˜2.2), which provides the ALS-based reference height maps, the dataset splits, and the vegetation mask. Then, we present the acquisition and preprocessing of the input Sentinel satellite time series (Section˜2.2.1).
@@ 2.2 Open-Canopy Dataset: Reference Data, Splits, and Vegetation Mask
To train and evaluate SERA-H, we used the Open-Canopy dataset (Fogel et al., 2025), a public benchmark specifically designed for high-resolution canopy height estimation. Derived from the French LiDAR HD campaigns acquired between 2021 and 2023222https://geoservices.ign.fr/lidarhd, the Open-Canopy dataset provides ALS-based canopy height maps at 1.5 m resolution which served as reference for height prediction. For the purpose of our study, we downsampled these reference ALS-based canopy height images to 2.5 m using max-pooling aggregation.
The dataset spans 87,383 tiles of $1\times 1$ km distributed across France, ensuring a wide diversity of forest types and topographies. We adhered to the fixed spatial partition defined by Open-Canopy: 66,339 km2 for training, 7,369 km2 for validation, and 13,675 km2 for testing (Fig.˜1), with a 1 km buffer systematically applied around test areas to prevent spatial leakage. Validation tiles are used to monitor performance and adjust model hyperparameters during training, and test tiles are used to perform a final evaluation of model performance.
Finally, to ensure that performance metrics focused on relevant areas, we restricted the evaluation to vegetation-covered zones using the vegetation mask provided by the Open-Canopy dataset. This mask was constructed by the union of a LiDAR-derived mask (height >1.5>1.5 m) and IGN forest polygons [^3] https://geoservices.ign.fr/bdforet. Note that this mask was exclusively applied during the validation and test phases for metric calculation and was not used during the training phase.
Figure 1: Spatial distribution of the Open-Canopy dataset. The map shows the distribution of the dataset across France (Fogel et al., 2025). The inset zoom on a region of this dataset shows the grid-based partition where each small grid square represents an individual 1 $\times$ 1 km tile. These tiles are divided between the training (blue), validation (purple), and test (red) tile sets. To prevent spatial data leakage, the test tiles are grouped together into larger blocks and a 1 km exclusion buffer is systematically applied around these aggregated test areas.
@@ 2.2.1 Satellite Data: Sentinel-1 and Sentinel-2
Unlike the Open-Canopy benchmark which relied on commercial SPOT-6/7 imagery, our model deployed input imagery from the Sentinel-1 and Sentinel-2 sensor systems, key components of the European Space Agency’s (ESA) Copernicus Earth observation program. These satellite systems provide complementary radar and optical data with global coverage and high temporal resolution.
For Sentinel-2, we utilized the Harmonized Sentinel-2 Level-2A Surface Reflectance product [^4] Available at: https://developers.google.com/earth-engine/datasets/catalog/COPERNICUS_S2_SR_HARMONIZED. We selected 10 spectral bands specifically for their sensitivity to forest structural and biochemical properties: the visible RGB channels (B2, B3, B4) provide information on canopy texture and shadows, the vegetation red-edge (B5, B6, and B7) and near-infrared (NIR, B8 and B8A) bands help capture leaf area index and vegetation density, and the short-wave infrared (SWIR, B11 and B12) bands offer insights into canopy moisture and biomass. Regarding Sentinel-1, we used the C-band Synthetic Aperture Radar (SAR) Ground Range Detected (GRD) product [^5] Available at: https://developers.google.com/earth-engine/datasets/catalog/COPERNICUS_S1_GRD and we extracted both VV and VH polarizations. While VH polarization is primarily sensitive to volume scattering within the branches and foliage, VV polarization provides complementary information on vertical structures and surface interactions. We further integrated both ascending (south-to-north) and descending (north-to-south) orbits to provide the model with opposite viewing angles (Fig.˜2b and Fig.˜2c). While providing additional information, this dual-orbit perspective helps minimize geometric distortions and radar shadows. Altogether, since these diverse spectral and backscatter signals from Sentinel-2 and Sentinel-1 are fundamentally correlated with tree height, the network learns to interpret their complex non-linear relationships to transform these multi-modal features into a canopy height prediction.
@@
Figure 2: Examples of input and reference data for a patch from the dataset. (a) Sentinel-2 image, (b) Sentinel-1 ascending image, and (c) Sentinel-1 descending image form one triplet of the time-series used as model input. (d) The corresponding reference is a canopy height map from the Open-Canopy dataset, resampled at 2.5 m. For visualization purposes, only a subset of the input bands is shown here (RGB for Sentinel-2 and $V$ for Sentinel-1). However, each triplet used by the model comprises 14 channels : 10 bands from $S2$, 2 from $S1_{\text{asc}}$ (VV/VH), and 2 from $S1_{\text{dsc}}$ (VV/VH).
We retrieved Sentinel images using the geefetch package [^6] https://geefetch.readthedocs.io/en/latest/ (Belouze et al., 2025), which interfaces with the Google Earth Engine (GEE) API [^7] https://earthengine.google.com/. For each tile of the Open-Canopy Dataset, we collected all Sentinel images available within a 120-day window centered on the acquisition date of the corresponding ALS data. To minimize the influence of atmospheric perturbations, Sentinel-2 data were subjected to a cloud-masking procedure. Images with more than 40% cloudy pixels or 20% high-probability cloud coverage were excluded. Individual pixels with a cloud probability above 40% or flagged in Sentinel-2’s QA60 band were masked.
Within this 120-day window, each Sentinel-2 ($S2$) image was paired with the temporally closest Sentinel-1 ascending ($S1_{\text{asc}}$) and descending ($S1_{\text{dsc}}$) acquisitions to form a triplet image $(S2,\allowbreak S1_{\text{asc}},\allowbreak S1_{\text{dsc}})$ (Fig.˜2). Therefore, each triplet image comprises 14 channels: 10 for Sentinel-2 bands, 2 for Sentinel-1 ascending (VV/VH), and 2 for Sentinel-1 descending (VV/VH). All these bands are harmonized at a resolution of 10 m using nearest-neighbor resampling. By forming all possible triplets from the available Sentinel images, we create a corresponding time-series of triplet images for each tile. Across the entire dataset, the average length of each time series is 13.51 triplets per tile of the Open-Canopy dataset. Apart from the processing described above, no additional filters were applied to the Sentinel images. The goal is to enable the model to learn how to filter out noise, such as atmospheric artifacts or speckle, by leveraging the temporal redundancy of the time series.
@@ 2.3 Model
Figure 3: Overview of the SERA-H model workflow. Each input sample consists of a temporal sequence of triplets, where each triplet is composed of one Sentinel-2 optical image and its temporally closest Sentinel-1 ascending and descending acquisitions $(S2,S1_{\text{asc}},S1_{\text{dsc}})$. This yields an input tensor of size $T\times C\times H\times W$, where $T$ is the fixed number of triplets in the temporal sequence, $C$ the total number of spectral and radar channels per triplet, and $H$ and $W$ the spatial dimensions of each image. The sequence is first processed by the EDSR super-resolution module, which upsamples all channels by $\times 4$ to match the reference ALS resolution. The super-resolved images are then passed through the UTAE encoder-decoder, which compresses the full temporal sequence using a temporal attention mechanism and outputs a single high-resolution canopy height map ($1\times 1\times 4H\times 4W$). Training is performed end-to-end using a Smooth L1 loss against ALS reference data, with gradients propagating through both UTAE and EDSR.
@@ 2.3.1 Enhanced Deep Super-resolution (EDSR) Module
Super-resolution aims to reconstruct a high-resolution image from low-resolution acquisitions by attempting to restore missing information. However, due to the very nature of this task, super-resolution algorithms are prone to creating hallucinations, i.e., non-existent details. These algorithms are systematically confronted with the perception-distortion trade-off (Blau and Michaeli, 2018): the mathematical impossibility of simultaneously optimizing digital fidelity (proximity to actual pixel values) and visual realism (perceived sharpness). With a view to producing accurate canopy height maps, we have prioritized measurement accuracy and the reduction of these artifacts. We have therefore selected the EDSR (Enhanced Deep Super-Resolution) architecture, which has demonstrated excellent reconstruction performance in terms of fidelity to real values (Blau and Michaeli, 2018; Karwowska and Wierzbicki, 2022; Ren et al., 2025).
Derived from the ResNet architecture, EDSR optimizes the model by removing unnecessary normalization layers. This simplification allows for a much larger and more powerful network to be built. We thus adopted this architecture in our workflow to upsample each individual image within the time series, increasing the spatial resolution by a factor of 4.
@@ 2.4.1 Training Data Sampling
During training, $1\times 1$ km tiles were randomly sampled from the Open-Canopy dataset. The corresponding time series of $(S2,S1_{\text{asc}},S1_{\text{dsc}})$ triplet images was retrieved. Tiles containing fewer than 4 distinct triplets were discarded. For the remaining tiles, as the model requires a fixed time series length of $T=16$, sequences shorter than this were padded by randomly duplicating existing triplets, while for longer sequences, we selected the $T=16$ triplets closest to the ALS acquisition date. The date assigned to each triplet corresponds to its Sentinel-2 acquisition date. Within the time series, triplets are then sorted by date. Finally, we cropped each triplet into $64\times 64$ pixel patches to match the model’s input dimensions. Corresponding ALS canopy height references were extracted at the same coordinates with a 2.5 m resolution, yielding $256\times 256$ pixel reference images.